# Beyond Sociodemographic Prompting: Using Supervision to Align LLMs with Human Response Distributions

### v1

@@ Abstract
The ability to accurately predict how different population groups would answer subjective questions would have great value. In this work, we show that use of relatively simple supervision can greatly improve language model alignment with diverse population groups, as measured over three datasets spanning various topics. Beyond evaluating average performance, we also report how alignment varies across specific groups. The simplicity and generality of our approach promotes easy adoption, while our broad findings provide useful guidance for when to use or not use our approach in practice. By conducting evaluation over many LLMs and prompting strategies, along with open-sourcing our work, we provide a useful benchmark to stimulate future research 111Our data and code will be available at https://github.com/GauriKambhatla/supervised-llm-alignment.
@@ 1 Introduction
Human surveys elicit valuable public opinion on a wide range of topics, such as proposed legislation, candidates for office (i.e., election polls), marketing or outreach campaigns, commercial products and services, etc. Hayati et al. (2024); Santurkar et al. (2023). Subjective annotation tasks in NLP also seek diverse, human judgments Sap et al. (2022); Biester et al. (2022); Pei and Jurgens (2023). Given the frequency, breadth, and importance of obtaining opinion data from diverse populations, there has been great interest in developing effective ways to conduct such surveys accurately and efficiently.
Recent work has sought to simulate human responses using LLMs in many areas, such as in psychology, sociology, and economic studies Aher et al. (2023), NLP annotation Nasution and Onan (2024), and large-scale survey creation and testing Rothschild et al. (2024). More specifically, LLMs are being increasingly used to simulate social or demographic groups Hu and Collier (2024). This is commonly done with persona or sociodemographic (SD) prompting Hu and Collier (2024); Beck et al. (2024) by both technical and non-technical practitioners alike Beck et al. (2024).
@@
Prior work on SD prompting has typically suffered from one or more key shortcomings. First, many studies have assumed a single majority answer for each sociodemographic group in modeling and/or evaluation Hwang et al. (2023); Hu and Collier (2024); Sun et al. (2025); Mukherjee et al. (2024). The fallacy in such a framing is clear: because members of a given group do not all share the same beliefs, accurate modeling and evaluation must incorporate intra-group disagreement. In addition, prompting for a distribution, rather than a representative response, could potentially make a model less susceptible to caricature and stereotyping Cheng et al. (2023a); Wang et al. (2024); Cheng et al. (2023b), better aligning with the goals of pluralistic alignment Sorensen et al. (2024). Prior studies modeling distributional beliefs have also tended to evaluate only a single method of extracting an LLM distribution Santurkar et al. (2023); Sun et al. (2024). Other work seeking to usefully replicate ground-truth answer distributions have been agnostic to underlying demographics, i.e., evaluating LLM distributions prompted with specific demographics with ground-truth distributions of multiple other demographics Beck et al. (2024).
@@
In this work, we evaluate SD prompting in a distributional manner across three large survey datasets (both US and global), using a variety of methods to extract distributions from LLMs. We focus on methods that work most broadly, supporting use with both black-box and white-box LLMs. We apply supervised calibration to better align LLM-generated distributions with human response distributions, with the intuition that LLM distributions might be directionally correct, but simply uncalibrated; e.g., LLM distributions might exaggerate differences between different groups.
@@
Concretely, we evaluate and explore distributional opinion alignment by studying the following research questions: (RQ 1) Does SD prompting generate distributions that are more aligned with human opinion? (RQ 2) Can we calibrate LLM-generated distributions to be more aligned with humans through supervised data? Is this consistent across models, datasets, and distribution elicitation methods? (RQ 3) Can generated distributions be more easily aligned with some SD groups over others? What is the effect of calibration for individual groups? (RQ 4) Does calibration work with fewer supervised training examples? And (RQ 5) What are the effects of post-training (methods like SFT and DPO) on LLM-generated distributions?
@@
Our experiments evaluate the alignment of LLM-generated distributions to human opinion, across three distribution elicitation methods, and 15 models of varying degrees of openness (open-source, open-model, API-access only), size, pre-training data, and post-training methods. We find that while baseline SD prompting techniques do not lead to consistent improvements in alignment across settings (RQ1), our approach to supervised calibration of LLM-generated distributions improves alignment by 16.3% on average across settings (RQ2). Moreover, as few as 5 gold examples can be used to calibrate distributions (RQ4), though alignment to some demographics is degraded more than others (RQ3). In regard to LLM post-training (RQ5), the effect appears to be dataset dependent, though we find that calibrated alignment is less affected than uncalibrated distributions.
We conclude with concrete suggestions to practitioners seeking to predict diverse human responses (e.g., public opinion surveys) via LLMs, offering appropriate caution.
@@ 2 Related Work
Sociodemographic (SD) Prompting
To support safe deployment of models, it is important that modeling and evaluation practices account for social dimensions of diversity Santurkar et al. (2023) that shape our social identities, including sociodemographic factors such as age, gender, and ethnicity Hwang et al. (2023); Movva et al. (2024). SD prompting refers to incorporating demographic variables into input prompts to enhance model alignment with desired outputs Alipour et al. (2024); Joshi et al. (2025). As past work has shown that neither annotator agreement methodologies nor reward mechanisms often produce significant differences in the solution Yang et al. (2024), prompting methods may support more efficient processes. Some work has reported that SD prompting can be effective in improving model performance Beck et al. (2024), supporting SD-specific diversification of predictions without need for collecting sensitive annotator SD information Gupta et al. (2023).
Pitfalls of Sociodemographic Prompting
@@
Prior studies evaluating sociodemographic prompting have reported mixed results in effectiveness in accurately simulating a social group Beck et al. (2024); Santurkar et al. (2023); Hu and Collier (2024). While some works show that sociodemographic prompting can improve performance, results vary greatly by prompt, model, and task Beck et al. (2024); Mukherjee et al. (2024).
@@
Other work warns against such methods due to misportrayal, othering, and exoticization of identities Cheng et al. (2023a); Wang et al. (2024); Cheng et al. (2023b). Simply incorporating demographic features into prompts does not always enable LLMs to adopt the perspectives of specific demographic groups Sun et al. (2025). Furthermore, a vast body of work has established biases in language models that may be perpetuated with sociodemographic prompting Nadeem et al. (2020); Gallegos et al. (2024); Gupta et al. (2024). There is risk of "pigeonholing particular demographic groups into specific narratives" Cheng et al. (2023a).
Distribution Elicitation
Confidence elicitation is the process of estimating LLM’s confidence in their responses without model fine-tuning or accessing internal information Geng et al. (2024a). Recent research has explored methods for confidence elicitation which are more suitable for both closed-source commercial APIs as well as open-source models Lin et al. (2024). Similarly, distribution elicitation methods are techniques used to obtain probability distributions from experts about unknown quantities Falconer et al. (2022).
@@ 3.1 Datasets
We consider three human survey datasets: Welcome Global Monitor 2018 (WGM), OpinionQA (OQA), and the World Values Survey (WVS). Both OQA and WVS have been used to study sociodemographic prompting in prior work Santurkar et al. (2023); Durmus et al. (2024). These three datasets cover a diverse range of subjective topics, such as perceptions of science and public health, public opinion on gun control, data privacy, and various moral opinions and values. To reduce downstream LLM costs, we subset datasets to include all WGM ordinal questions and two questions per category from OQA and WVS. This totals 92 questions across WGM (14), OQA (38), and WVS (40).
Demographics for each dataset are shown in Table 11 (Appendix). Over all datasets, questions, and demographics, there are 4,500 human response distributions. For each question, model, and distribution elicitation method (Section 3.2), we predict the probability distribution of human responses over answer choices. This yields 220,500 generated SD-specific response distributions and 4,500 generated SD-agnostic response distributions.
@@ 3.2 Distribution elicitation
From ground truth human response data, we infer reference probability distributions over answer choices for each question by relative frequency, specific to each SD group. To predict these distributions via LLM, we apply (with some modifications) methods inspired from confidence literature as follows. We investigate techniques that work most broadly, supporting use with both black-box and white-box LLMs. That said, Appendix A.3 reports a smaller scale study with log probabilities. Appendix B shows prompts used for all methods.
Verbalized. This is simply directly prompting the model to output a distribution in numbers Geng et al. (2024b); Tian et al. (2023). For example, given the question How much do you trust vaccines?, where the answer choices are a lot, somewhat, not much, and not at all, we prompt the model to output a distribution over the answer choices, such as [0.7, 0.2, 0.05, 0.05]. We sample times and average for variance reduction. If a model ever does not generate a distribution as requested, we simply discard that output. We also renormalize generated distributions that do not sum to 1.
Self-random. We prompt the model to output a single answer choice, sampling times Xiong et al. (2024), with a temperature of 0.7. We then create a distribution over the responses. For example, given the question How much do you trust vaccines?, where the answer choices are a lot, somewhat, not much, and not at all, we prompt the model to output a single number that correlates with an answer choice, such as 4 for a lot or 1 for not at all. Note that this method estimates the log probabilities for large Tian et al. (2023). In our experiments, we sample times.
@@ 3.3 Prompts
We use two categories of prompts, and for each category, we prompt with the different elicitation methods. See Appendix B for prompts used.
@@
Sociodemographic (SD). These use a form similar to “Imagine you are {d}. Question: ”, where {d} is a demographic value. The exact prompt varies with demographic and elicitation method. See Appendix Table 13 for full prompts used.
3.4 Metric
We use the opinion alignment metric from Santurkar et al. (2023) to measure similarity between elicited distributions and ground truth distributions for all results. This metric makes use of 1 Wasserstein distance, or earth-mover’s distance. As noted by Santurkar et al. (2023), this metric takes into account the ordinal nature of the survey questions, as opposed to other distribution divergence metrics like Kullback-Liebler or Jensen-Shannon. The opinion alignment metric ranges from 0 to 1, but in our results we show the value as a percentage.
@@ 3.5 Models
3.6 Calibration
We apply supervised regression to transform LLM-generated distributions. We split the LLM distributions into train, development, and test sets (60-20-20). Results in Section 6 show that far less supervision is typically needed in practice.
The input to regression is the LLM distribution and the output is the scaled distribution. We split the distributions by answer choice before normalizing across choices. More formally, for a regression model and LLM distribution parameterized by the values for each answer choice , we learn a regression such that each value is transformed using supervision from ground truth values for each answer choice . I.e., we train on pairs of to learn transformed values on held-out test questions. We learn a regression model for each dataset-LLM-elicitation setting and do development set hyperparameter tuning and model selection to choose between linear regression (including lasso and ridge) and random forest regression.
@@ 4 Aligning LLMs with human opinions
We compare two methods: adding SD information to prompts, and calibrating LLM-generated distributions using some human ground truth data for supervision. We evaluate the effectiveness of these two methods in aligning with opinion distributions of those SD groups across models, datasets, and probability elicitation methods. Our main results are shown in Table 1. The full results for all models are shown in Appendix Table 5.
Findings (RQ1, RQ2):
(1) the effectiveness of baseline SD prompting in generating aligned distributions is model, dataset, and elicitation method dependent, aligning with prior work that studies this in the non-distributional setting Beck et al. (2024); (2) calibration increases opinion alignment in aggregate across models, datasets, and elicitation methods by an average of 16.3%; and (3) calibration reduces variance within and across settings.

### v2

# Improving the Distributional Alignment of LLMs using Supervision

@@ Abstract
The ability to accurately align LLMs with human population groups on subjective questions would have great value. In this work, we show that use of simple supervision can greatly improve language model alignment with diverse population groups more consistently, as measured over three datasets spanning various topics. Beyond evaluating average alignment, we also report how alignment varies across specific groups. Our broad findings provide insights into the distributional alignment of LLMs with diverse population groups. By conducting evaluation over many LLMs and prompting strategies, along with open-sourcing our work, we provide a benchmark to stimulate future research.111Our data and code will be available at https://github.com/GauriKambhatla/supervised-llm-alignment
@@ 1 Introduction
As LLMs are increasingly incorporated into real-world systems, it is more and more important to evaluate their alignment with human values and opinions. A growing body of work seeks to align LLMs with particular human population groups, with the goal of simulating their opinions or behaviors, by steering LLM generations towards those groups Aher et al. (2023); Sun et al. (2024); Dillion et al. (2023). One common method to do so is persona or sociodemographic (SD) prompting, which refers to defining a persona with demographic or description variables within a prompt, with the goal of modeling the behavior of that group using an LLM Hu and Collier (2024); Beck et al. (2024).
@@
Recent work in NLP has sought to evaluate SD prompting by studying its performance on large surveys and polls, which contain many subjective questions, along with hundreds of thousands of human responses Santurkar et al. (2023); Durmus et al. (2024); Masoud et al. (2025); AlKhamissi et al. (2024). However, evaluation of SD prompting has typically suffered from one or more key shortcomings. First, many studies have assumed a single majority answer to each question for each sociodemographic group in modeling and/or evaluation Hwang et al. (2023); Hu and Collier (2024); Sun et al. (2025a); Mukherjee et al. (2024); Durmus et al. (2024); Masoud et al. (2025); AlKhamissi et al. (2024). The fallacy in such a framing is clear: because members of a given group do not all share the same beliefs, accurate modeling and evaluation must incorporate intra-group disagreement. Prompting for a distribution, rather than a representative response, could potentially also make a model less susceptible to caricature and stereotyping Cheng et al. (2023a); Wang et al. (2025a); Cheng et al. (2023b), better aligning with the goals of pluralistic alignment Sorensen et al. (2024). However, Prior studies modeling distributional beliefs have also tended to evaluate only a single method of extracting an LLM distribution Santurkar et al. (2023); Sun et al. (2024), or evaluate LLM distributions prompted for specific demographics in a demographic-agnostic manner Beck et al. (2024).
@@
In this work, we evaluate SD prompting in a distributional manner across three survey datasets using a variety of methods to extract distributions from LLMs. We focus on methods that work most broadly, supporting use with open-source, open-model, and API-access LLMs. We apply supervised calibration to better align LLM-generated distributions with human response distributions, with the intuition that LLM distributions might be directionally correct, but simply uncalibrated; e.g., LLM distributions might exaggerate differences between different groups. We use supervised calibration to mean learning a transformation to a distribution to be more similar to human response distributions, using human responses as supervised labels (described in detail in Section 3.6). We use alignment to mean the degree to which opinions align between humans and LLMs (see Section 3.4).
@@
Concretely, we investigate distributional opinion alignment via the following research questions: (RQ 1) Does SD prompting generate distributions that are more aligned with human opinion? (RQ 2) Can we calibrate LLM-generated distributions to be more aligned with humans through supervised data? Is this consistent across models, datasets, and distribution elicitation methods? (RQ 3) Can generated distributions be more easily aligned with some SD groups over others? What is the effect of calibration for individual groups? (RQ 4) Does calibration work with fewer supervised training examples? And (RQ 5) What are the effects of post-training (doing fine-tuning and reinforcement learning) on LLM-generated distributions?
@@
We evaluate the alignment of LLM-generated distributions to human opinion across 3 distribution elicitation methods and 15 models of varying degrees of openness (open-source, open-model, API-access only), size, pre-training data, and post-training methods. We first show that baseline SD prompting techniques do not lead to consistent improvements in alignment across settings (RQ1). In contrast, our approach to supervised calibration of LLM-generated distributions improves alignment by 16.3% on average across prompt methods, elicitation methods, datasets, and LLMs (RQ2), though it improves for some SD groups and degrades for others (RQ3). Moreover, as few as 1-10 gold examples can be used to calibrate distributions (RQ4). In regard to LLM post-training (RQ5), calibration results in more consistent behavior across various post-training methods, both with and without SD prompting; without calibration, alignment is much more dependent on both the dataset and the post-training method used.
We conclude with a discussion of the methods and models that generate the distributions most aligned with human groups on subjective questions. Caution is warranted given varying success in aligning to different groups as well as recognition that such efforts risk essentializing sociodemographic categories if not applied with care.
@@ 2 Related Work
Social alignment of LLMs
LLMs are increasingly used to simulate social human groups in a variety of settings, such as psychology, sociology, and economic studies Aher et al. (2023); Dillion et al. (2023), survey development Rothschild et al. (2024); Jansen et al. (2023), NLP annotation Nasution and Onan (2024); Hashemi et al. (2024); Movva et al. (2024), and multi-agent interactive simulation Hou et al. (2025); Zhou et al. (2024).
Many methods have been proposed to better align LLMs with human values and opinions Ma et al. (2024). For example, persona or sociodemographic (SD) prompting defines a persona with demographic or description variables within a prompt to simulate or model groups Hu and Collier (2024); Beck et al. (2024). Other methods include: using knowledge of a user’s previous opinions to model their future opinions Hwang et al. (2023); Gordon et al. (2022), using psychological scaffolds to improve model rationalization Joshi et al. (2025), self-alignment via the simulation of social scenes Pang et al. (2024), extracting persona vectors to mitigate certain character traits Chen et al. (2025), creating more descriptive value profiles Sorensen et al. (2025), using in-context learning Choenni and Shutova (2025), and fine-tuning LLMs followed by collection of a small amount of human data Krsteski et al. (2025). Evaluation frameworks have also been proposed for assessing alignment and social simulation Kang et al. (2025); Shi et al. (2025); Lutz et al. (2025).
Pitfalls of social alignment
@@
Various limitations have been reported, such as LLMs being unreliable as personalized annotators Dong et al. (2024), unable to use different languages to steer generation towards groups Kwok et al. (2024), biased towards some groups Sun et al. (2025b), lacking stability and steerability under prompt variations Khan et al. (2025), and lacking the reasoning depth to simulate humans in social experiments Gao et al. (2025). Prior work evaluating SD prompting has also reported mixed results in accurately simulating social groups Beck et al. (2024); Santurkar et al. (2023); Hu and Collier (2024); Kaiser et al. (2025). While some works show that SD prompting can improve performance, results vary greatly by prompt, model, and task Beck et al. (2024); Mukherjee et al. (2024); Bisbee et al. (2024).
@@
Other work warns against such methods due to misportrayal, othering, and exoticization of identities Cheng et al. (2023a); Wang et al. (2025a); Cheng et al. (2023b); Kitadai et al. (2024). Simply incorporating demographic features into prompts does not always enable LLMs to adopt the desired perspectives of specified demographic groups Sun et al. (2025a). Furthermore, a vast body of work has established biases in language models that may be perpetuated with SD prompting Nadeem et al. (2021); Gallegos et al. (2024); Gupta et al. (2024).
Probabilistic reasoning with LLMs
Prior work in NLP has studied the use of LLMs for eliciting and working with uncertainty and probability distributions. Paruchuri et al. (2024) study whether LLMs can reason about probability distributions, finding that LLMs are able to answer questions about distributions, though performance varies based on the type of distribution. Wang et al. (2025b) develop a model and framework for modeling and evaluating fine-grained probability estimation of propositions conditioned on context, and show that their model captures some amount of human uncertainty. Prior work has also found that several LLMs are capable few-shot regressors, often outperforming traditional supervised methods Vacareanu et al. (2024), and capable fine-tuned regressors Song et al. (2024).
Distribution prediction in the real world is a complex task, and can rely on both uncertainty quantification and nuanced language understanding Wang et al. (2025b); Vedula et al. (2025). This motivates using LLMs for this task, and prior work has developed methods in various domains for doing LLM-based distribution prediction: combining LLM-generated distributions to better predict human annotator distributions Hashemi et al. (2024), developing a quantile regression approach for price prediction Vedula et al. (2025), and mapping LLM token predictions to continuous distributions for probabilistic forecasting Gruver et al. (2023).
Distributional alignment
Recent works have also advocated for a shift to pluralistic alignment due to the ecological fallacy of assuming all members of a group share the same values, beliefs, or opinions Gordon et al. (2022); Sorensen et al. (2024); Feng et al. (2024); Kirk et al. (2024); Xie et al. (2025). Closer to our work, prior works have studied aligning LLM distributions with human response distributions on subjective questions. We draw inspiration from Santurkar et al. (2023), though we evaluate multiple methods of extracting distributions from LLMs. Dominguez-Olmedo et al. (2024) study systematic biases with LLM-generated distributions and how this differs from human response distributions for surveys. Meister et al. (2025) study different distribution elicitation methods for subjective questions. Cao et al. (2025) fine-tune with a divergence-based loss function. Our work introduces the notion of supervised calibration in this space, and studies the effect of alignment methods on individual groups.
@@ 3.1 Datasets
We consider three survey datasets: Welcome Global Monitor 2018 (WGM), OpinionQA (OQA) Santurkar et al. (2023), and the World Values Survey (WVS). OQA and WVS have been used to study alignment in prior work Durmus et al. (2024); Masoud et al. (2025); AlKhamissi et al. (2024). These three datasets cover diverse subjective topics, including global perceptions of science and public health (WGM), US public opinion (OQA), and various moral opinions and values (WVS). We include all WGM ordinal questions and two questions per category from OQA and WVS, yielding 92 questions across WGM (14), OQA (38), and WVS (40) and a total of 4,500 human response distributions over all datasets, questions, and demographics.
@@ 3.2 Distribution elicitation
From ground truth human response data, we infer reference probability distributions over answer choices for each question by relative frequency, specific to each SD group. To generate distributions via LLM, we apply (with some modifications) methods inspired from model confidence literature. We investigate techniques that work most broadly, supporting use with open-source, open-model, and API-access LLMs. That said, Appendix A.2 reports a smaller scale study with log probabilities. Appendix B shows prompts used for all methods.
Verbalized. This is directly prompting the model to output a probability distribution Geng et al. (2024); Tian et al. (2023). For example, given the question How much do you trust vaccines?, where the answer choices are a lot, somewhat, not much, and not at all, we prompt the model to output a distribution over the answer choices, such as [0.7, 0.2, 0.05, 0.05]. We sample times and average for variance reduction. If a model ever does not generate a distribution as requested, we simply discard that output. We also renormalize generated distributions that do not sum to 1.
Self-random. We prompt the model to output a single answer choice, sampling times Xiong et al. (2024), with a temperature of 0.7. We then create a distribution over the responses. For example, given the question How much do you trust vaccines?, where the answer choices are a lot, somewhat, not much, and not at all, we prompt the model to output a single number that correlates with an answer choice, such as 4 for a lot or 1 for not at all. Note that this method estimates the log probabilities for large Tian et al. (2023). In our experiments, we sample times, following prior work Xiong et al. (2024).
@@ 3.3 Prompts
We use two categories of prompts, with three variations. For each, we prompt with the 3 different elicitation methods. See Appendix B for the prompts.
@@
Sociodemographic (SD). These use a form similar to “Imagine you are {d}. Question: ”, where d is a demographic value. The exact prompt varies with demographic and elicitation method. See Appendix Table 15 for full prompts.
3.4 Metric: Opinion Alignment
We use the following opinion alignment metric from Santurkar et al. (2023) to measure similarity between an elicited distribution () and ground truth distribution () for the set of questions :
where refers to Wasserstein distance, or earth-mover’s distance, and is a normalization factor. Santurkar et al. (2023) note this metric accounts for the ordinal nature of the survey questions, as opposed to other distribution divergence metrics like Kullback-Liebler or Jensen-Shannon. The opinion alignment metric ranges in , but in our results we show the value as a percentage.
@@ 3.5 Models
3.6 Calibration
We apply supervised regression to transform LLM-generated distributions to more closely align them with the ground truth human distributions. For example, for the question, How much do you trust doctors and nurses? A lot, some, not much, or not at all?, the LLM might generate a distribution such as , while the ground truth human response distribution based on the survey question might be . We seek to learn a transformation from the LLM-generated distributions to the human distributions that might be applied to all LLM-generated distributions. More formally, for a regression model and LLM distribution parameterized by the values for each answer choice (where is the number of answer choices), we learn a regression such that each value is transformed using supervision from ground truth values for each answer choice . We then re-normalize the transformed distribution.
Training examples. The examples for the regression model are the distributions split by answer choices. For example, the LLM and human distributions given above, and , would be split into supervised pairs of , , , and . We thus train on pairs of to learn transformed values on held-out test questions.
Regression model. We learn a regression model for each dataset-LLM-elicitation-prompt setting (270 regression models for 3 datasets, 15 models, 3 elicitation methods, and 2 prompt categories). Hyperparameter tuning and model selection are used to choose between linear regression (including lasso and ridge) and random forest regression for each regressor. We use the scikit-learn Pedregosa et al. (2011) implementation for each regressor.
Training. The total number of examples for each regression model is x x , where is the class within each sociodemographic group (e.g., female and male for sex), is the question (e.g., how much do you trust doctors?), and is the number of answer choices per question. These values are different for each setting (see Table 13 for the sizes of and per dataset). We were unable to extract distributions for all questions from some of the smaller LLMs (see Table 17), so the number of training examples differs for them. We split the LLM distributions into train, development, and test sets (80%-10%-10%).
For minimal supervision (Section 6), we train on a random sample of examples ( = [1, 5, 10, 50, 100, 200, 500, 1000, 1200], where each example is the full distribution not split by answer choices) from the training set. We repeat this procedure 10 times for different random splits and average.
Evaluation. We optimize the regression models with Mean Squared Error (MSE). Since we split distributions during training, we normalize output values so that they sum to 1 before calculating opinion alignment with the ground truth distribution. For example, if the output of for the distribution is , we would normalize these values to .
@@ 4 Aligning LLMs with human opinions
We compare two methods: adding SD information to prompts and calibrating LLM-generated distributions with human ground truth data for supervision. We evaluate the effectiveness of these two methods in aligning with opinion distributions of those SD groups across models, datasets, and probability elicitation methods. Our main results are shown in Table 1. The alignment values for all models are shown in Appendix Table 7 for all the results in this section. All significance results are calculated via paired t-test with Bonferroni correction.